# A Correspondence-Driven Approach for Bilevel Decision-making with Nonconvex Lower-Level Problems

### v1

@@ 3.1 Definition of the Correspondence-Driven Hyperfunction
Instead of modeling the follower as a rational agent who always attains a global minimizer, we consider the follower to be an algorithm-based bounded rational agent. Specifically, the follower’s decision is generated by executing a prescribed optimization method $\mathcal{M}$ with step size schedule $\bm{\eta}$ from a fixed initialization $y_{0}$. For any given upper-level variable $x$, this procedure yields a sequence, and $\hat{S}(x,y_{0},\bm{\eta},\mathcal{M})$ denotes the subset of its accumulation points attaining the smallest lower-level objective value, from which the follower selects its decision. Regarding the algorithm $\mathcal{M}$ and step size schedule $\bm{\eta}$, we impose the following mild assumption, which can be satisfied by a wide range of standard optimization methods, such as the gradient descent method.
@@
The accumulation points of the sequence generated using method $\mathcal{M}$ with step size schedule $\bm{\eta}$ for initial point $y_{0}$ and any upper-level decision variable $x$ are stationary points with respect to $y$. That is, the set of all limit points of the sequence is a subset of the lower-level stationary point set.
@@
Given an upper-level decision variable $x\in\mathcal{X}$, a fixed follower initialization $y_{0}\in\mathcal{Y}$, a step size schedule $\bm{\eta}=\{\eta_{t}\}$, and an optimization method $\mathcal{M}$ satisfying Assumption 3.1, let $y^{\mathrm{cd}}(x,y_{0},\bm{\eta},\mathcal{M})$ defined as follows:
@@
This extension captures the average behavior of the follower under randomized initializations and reduces to (3.2) when $\mathcal{Y}_{0}=\{y_{0}\}$. Our proposed approach SCiNBiO can be extended to this setting as well; however, for simplicity of analysis, we focus on the single-initialization case in this paper.
@@ 3.2 Smoothing via Mollification
whose graph shifts left or right as $x$ varies. Figure 3 illustrates the function for $x=0$. We fix the initialization $y_{0}=0$ and apply gradient descent with step size smaller than $1/L$, where $L$ is the Lipschitz constant of $\nabla_{y}g(x,y)$ in the compact sublevel set $\{y:g(x,y)\leq g(x,0)\}$. In this setting, the algorithm converges to different local minimizers depending on the value of $x$: for $x<0$, the trajectory converges to the right minimizer; for $x>0$, it converges to the left minimizer; and for $x=0$, the algorithm stagnates at the maximizer $y^{\text{cd}}(0)=0$. As a result, $y^{\text{cd}}(x)$ is discontinuous at $x=0$, and consequently, $\phi^{\text{cd}}=f(x,y^{\text{cd}}(x))$ is also discontinuous at that point unless the upper-level objective $f$ is carefully designed to cancel the jump.
@@
This smoothing procedure produces a smooth approximation $\phi^{\text{cd}}_{\xi}(x)$, which approximates $\phi^{\text{cd}}(x)$ well in regions where $\phi^{\text{cd}}(x)$ is continuous. As illustrated in Figure 4, the smoothed function aligns well with the original function except near the discontinuity. Moreover, the minimizer of the smoothed function converges to $0^{-}$ as the smoothing parameter $\xi$ tends 0.
@@
  • The second issue concerns the integrability of $\phi^{\text{cd}}(x)$: since this function is generally discontinuous, it is not immediately clear whether the convolution integral and its gradient are well defined. In particular, the integrability of discontinuous functions requires careful analysis of the set of discontinuity points.
To ensure that the convolution $\phi^{\text{cd}}_{\xi}(x)$ is well defined and to provide a foundation for the theoretical developments that follow, we now introduce a set of assumptions.
@@
Assumption 3.2(1)-(5) are standard assumptions in the BLO literature. Assumption 3.2(6) is generally assumed in much of single-level optimization literature. Then we can directly obtain the following bounds by Assumptions 3.2(2)(6), which state the compactness of $\mathcal{X}$ and the uniform compactness of the level sets $\{y:g(x,y)\leq g(x,y_{0})\}$, respectively.
@@
    $f(x,y)$ is Lipschitz continuous with constant $L_{f}$ on $\mathcal{X}\times\{y:g(x,y)\leq g(x,y_{0})\}$;
  3. 3.
@@
Assumption 3.3 is mild and holds for a wide range of commonly used optimization methods (e.g., gradient descent, accelerated methods, or second-order methods) with properly selected step sizes. The relation (3.7) implies that the output $y^{\text{cd}}(x)$ lies in the sublevel set $\{y:g(x,y)\leq g_{0}^{\ast}\}$. Combining this with Proposition 3.1(3), which ensures that $f(x,y)\leq\overline{f}$ for all $y$ in this sublevel set, we obtain the uniform bound
@@
To address the second issue, namely the integrability of $\phi^{\mathrm{cd}}(x)$, we partition $\mathcal{X}$ into the set of bifurcation points $\widetilde{\mathcal{X}}$ defined in (2.3) and its complement $\mathcal{X}\setminus\widetilde{\mathcal{X}}$, and impose mild assumptions on each set separately. Leveraging the result that the discontinuity points of $y^{\mathrm{cd}}(x)$ within both sets have measure zero, we show that the integral is well defined. This is motivated by a classical result in real analysis: the integral of a bounded function is well defined as a Riemann integral if and only if its set of discontinuities has Lebesgue measure zero.
For the set $\mathcal{X}$, we use Assumption 2.1, which implies that the bifurcation point set $\widetilde{\mathcal{X}}$ defined in (2.3) has measure zero. For the remaining set $\mathcal{X}\setminus\widetilde{\mathcal{X}}$, we make the following assumption:
@@
Under Assumption 2.1 and 3.4, $y^{\text{cd}}(x)$ is discontinuous only on a set of measure zero in $\mathcal{X}$. Since $\phi^{\text{cd}}(x)=f(x,y^{\text{cd}}(x))$ and the function $f$ is continuous, it follows that $\phi^{\text{cd}}(x)$ is also discontinuous only on a measure-zero subset of $\mathcal{X}$. Note that $\mathcal{X}$ is a closed convex set, and thus its boundary $\partial\mathcal{X}$ has Lebesgue measure zero (Lang, 1986, Theorem 1). The extension of $\phi^{\text{cd}}(x)$, i.e., setting $\phi^{\text{cd}}(x)=\overline{f}$ for $x\notin\mathcal{X}$, introduces new discontinuities only on a subset of $\partial\mathcal{X}$, since the function takes different values inside and outside $\mathcal{X}$. Therefore, the set of discontinuities of the extended $\phi^{\text{cd}}(x)$ on $\mathbb{R}^{n}$ remains a measure-zero set. By the classical criterion for Riemann integrability, it follows that $\phi^{\text{cd}}(x)$ is Riemann integrable on $\mathcal{X}$. Therefore, the second technical issue is resolved. In conclusion, the convolution $\phi^{\text{cd}}_{\xi}(x)$ is well defined.
@@
Suppose Assumption 3.2 holds. Then $\phi^{\text{cd}}_{\xi}(x)$ is smooth for any $\xi$. Furthermore, if $\phi^{\text{cd}}(x)$ is continuous at point $\bar{x}$, then we have the convergence of function value $\lim_{\xi\to 0}\phi^{\text{cd}}_{\xi}(\bar{x})=\phi^{\text{cd}}(\bar{x})$. If $\phi^{\text{cd}}(x)$ is differentiable at point $\bar{x}$, then we have the convergence of gradient $\lim_{\xi\to 0}\nabla_{x}\phi^{\text{cd}}_{\xi}(\bar{x})=\nabla_{x}\phi^{\text{cd}}(\bar{x})$, and thus $\lim_{\xi\to 0}P_{\mathcal{X}}(x,\nabla_{x}\phi^{\text{cd}}_{\xi},\beta)=P_{\mathcal{X}}(x,\nabla_{x}\phi^{\text{cd}},\beta)$ which are defined as follows:
@@
Proposition 3.2 shows that Gaussian smoothing gives $\phi^{\text{cd}}_{\xi}(x)$ good regularity: it is smooth for all $\xi>0$, and it recovers both the function value and (proximal) gradient of $\phi^{\text{cd}}$ in the limit as $\xi$ approaches 0, at any point where $\phi^{\text{cd}}$ is continuous or differentiable. This result ensures that the projected gradient-based methods applied to the smoothed problem can approximate the behavior of the original hyperfunction. The proof follows the classical theory of mollification in PDE (Evans, 2022), where a discontinuous or nonsmooth function is approximated by the convolution with a smooth kernel. In classical mollifier theory, a compactly supported smooth bump function (e.g., supported on the unit ball) is often used. Pointwise and gradient convergence are established using local continuity or differentiability inside the support, while the contribution from outside the support is exactly zero due to compactness. In our case, the smoothing kernel is the Gaussian $h_{\xi}$, which is not compactly supported. Nevertheless, the key idea remains the same: most of the kernel mass is concentrated around the origin, and the contribution from the tail decays exponentially fast. This exponential decay plays the same role as compact support in standard mollifier arguments. Thus, the structure of the proof is essentially identical to the standard bump function case.
@@ 4.1 Sketch of the Proposed Algorithm
$=\frac{1}{\xi}\mathbb{E}_{u\sim\mathcal{N}(0,I^{n})}[uf(x+\xi u,y^{\text{cd}}(x+\xi u))].$
Note that, as established in (3.9), we have extended $\phi^{\text{cd}}$ outside $\mathcal{X}$, which is equivalently interpreted as redefining $f$ to take the constant value $\overline{f}$ outside $\mathcal{X}$, where $\overline{f}$ is from Proposition 3.1(3). Thus, the function $f$ should be understood in the sense of this redefinition.
In practice, we estimate this gradient via Monte Carlo sampling, which leads to a biased stochastic gradient descent (SGD) algorithm. Specifically, we draw i.i.d. sample points $u^{(1)},...,u^{(N)}\sim\mathcal{N}(0,I^{n})$, and for each such sample $u^{(i)}$, if the perturbed point $x+\xi u^{(i)}\notin\mathcal{X}$, we directly assign the value $f(x+\xi u^{(i)},\hat{y}(x+\xi u^{(i)}))=\overline{f}$; otherwise, when $x+\xi u^{(i)}\in\mathcal{X}$, we run a fixed number $K$ of iterations of the lower-level algorithm $\mathcal{M}$ at $x+\xi u^{(i)}$ to obtain an approximate solution $\hat{y}(x+\xi u^{(i)})$, where $K$ is independent of $x$. We then form the following average
@@
         Run inner algorithm $\mathcal{M}$ on $g(\tilde{x}^{(i)},y)$ for $K_{t}$ steps, initialized at $y_{0}$, with step size $\eta$ to obtain $\hat{y}^{(i)}$;
@@
  Compute the gradient estimator $\hat{\nabla}_{x}^{K_{t}}\phi^{\text{cd}}_{\xi}(x_{t})$ given by (4.3) and update $x_{t+1}$$\hat{\nabla}_{x}^{K_{t}}\phi^{\text{cd}}_{\xi}(x_{t})\leftarrow\frac{1}{N_{t}\,\xi}\sum_{i=1}^{N_{t}}u^{(i)}f^{(i)}$;
@@
$\nabla_{x}\phi^{\text{cd}}_{\xi}(x)-\nabla^{K}_{x}\phi^{\text{cd}}_{\xi}(x)=\int_{\mathbb{R}^{n}}\nabla h_{\xi}(x-z)[f(z,y^{\text{cd}}(z))-f(z,\hat{y}(z))]dz.$ (4.5)
@@
$\int_{\widetilde{\mathcal{X}}_{\delta}}\nabla h_{\xi}(x-z)[f(z,y^{\text{cd}}(z))-f(z,\hat{y}(z))]\,dz+\int_{\mathbb{R}^{n}\setminus\widetilde{\mathcal{X}}_{\delta}}\nabla h_{\xi}(x-z)[f(z,y^{\text{cd}}(z))-f(z,\hat{y}(z))]\,dz.$ (4.7)
@@
Finally, to obtain the lower-level problem oracle complexity of SCiNBiO, we introduce the notion of fold bifurcation Section 4.4 to endow the bifurcation points with additional structure. This refined characterization also gives us more geometric information of the bifurcation points set.

### v3

@@ 3.1 Definition of the Correspondence-Driven Hyperfunction
Instead of modeling the follower as a rational agent who always attains a global minimizer, we consider the follower to be an algorithm-based bounded rational agent. Specifically, the follower’s decision is generated by executing a prescribed optimization method $\mathcal{M}$ with step size schedule $\bm{\eta}$ from a fixed initialization $y_{0}$. For any given upper-level variable $x$, this procedure yields a sequence, and $\hat{S}(x,y_{0},\bm{\eta},\mathcal{M})$ denotes its accumulation points, from which the follower selects its decision. Regarding the algorithm $\mathcal{M}$ and step size schedule $\bm{\eta}$, we impose the following mild assumption, which can be satisfied by a wide range of standard optimization methods, such as the gradient descent method.
@@
The accumulation points of the sequence generated using method $\mathcal{M}$ with step size schedule $\bm{\eta}$ for initial point $y_{0}$ and any upper-level decision variable $x$ are stationary points with respect to $y$. That is, the set of all limit points of the sequence are a subset of the lower-level stationary point set.
@@
Given an upper-level decision variable $x\in\mathcal{X}$, a fixed follower initialization $y_{0}\in\mathcal{Y}$, a step size schedule $\bm{\eta}=\{\eta_{t}\}$, and an optimization method $\mathcal{M}$ satisfying Assumption 3.1, let $y^{\mathrm{cd}}(x,y_{0},\bm{\eta},\mathcal{M})$ be defined as follows:
@@
This extension captures the average behavior of the follower under randomized initializations and reduces to (3.2) when $\mathcal{Y}_{0}=\{y_{0}\}$. Our proposed approach SCiNBiO (Algorithm 1) can be extended to this setting as well; however, for simplicity of analysis, we focus on the single-initialization case in this paper.
@@ 3.2 Smoothing via Mollification
whose graph shifts left or right as $x$ varies. Figure 3 illustrates the function for $x=0$. We fix the initialization $y_{0}=0$ and apply gradient descent with step size smaller than $1/L$, where $L$ is the Lipschitz constant of $\nabla_{y}g(x,y)$ in the compact sublevel set $\{y:g(x,y)\leq g(x,0)\}$. In this setting, the algorithm converges to different local minimizers depending on the value of $x$: for $x<0$, the trajectory converges to the right minimizer; for $x>0$, it converges to the left minimizer; and for $x=0$, the algorithm stagnates at the maximizer $y^{\text{cd}}(0)=0$. As a result, $y^{\text{cd}}(x)$ is discontinuous at $x=0$, and consequently, $\phi^{\text{cd}}(x)=f(x,y^{\text{cd}}(x))$ is also discontinuous at that point unless the upper-level objective $f$ is carefully designed to cancel the jump.
@@
This smoothing procedure produces a smooth approximation $\phi^{\text{cd}}_{\xi}(x)$, which approximates $\phi^{\text{cd}}(x)$ well in regions where $\phi^{\text{cd}}(x)$ is continuous. As illustrated in Figure 4, the smoothed function aligns well with the original function except near the discontinuity. Moreover, the minimizer of the smoothed function converges to 0 from the left as the smoothing parameter $\xi$ tends to 0.
@@
  • The second issue concerns the integrability of $\phi^{\text{cd}}(x)$: since this function is generally discontinuous, it is not immediately clear whether the convolution integral and its gradient are well-defined. In particular, the integrability of discontinuous functions requires careful analysis of the set of discontinuity points.
To ensure that the convolution $\phi^{\text{cd}}_{\xi}(x)$ is well-defined and to provide a foundation for the theoretical developments that follow, we now introduce a set of assumptions.
@@
  7. 7.
    $f(x,y)$ is Lipschitz continuous with constant $L_{f}$ on $\mathcal{X}\times\{y:g(x,y)\leq g(x,y_{0})\}$.
Assumptions 3.2(1)-(5) and (7) are standard assumptions in the BLO literature. Assumption 3.2(6) is generally assumed in much of single-level optimization literature. Then we can directly obtain the following bounds by Assumptions 3.2(2)(6), which state the compactness of $\mathcal{X}$ and the uniform compactness of the level sets $\{y:g(x,y)\leq g(x,y_{0})\}$, respectively.
@@
Assumption 3.3 is mild and holds for a wide range of commonly used optimization methods (e.g., gradient descent, accelerated methods, or second-order methods) with properly selected step sizes. The relation (3.7) implies that the output $y^{\text{cd}}(x)$ lies in the sublevel set $\{y:g(x,y)\leq g(x,y_{0}))\}$. Combining this with Proposition 3.1(2), which ensures that $f(x,y)\leq\overline{f}$ for all $y$ in this sublevel set, we obtain the uniform bound
@@
To address the second issue, namely the integrability of $\phi^{\mathrm{cd}}(x)$, we partition $\mathcal{X}$ into the set of bifurcation points $\widetilde{\mathcal{X}}$ defined in (2.3) and its complement $\mathcal{X}\setminus\widetilde{\mathcal{X}}$, and impose mild assumptions on each set separately. Leveraging the result that the discontinuity points of $y^{\mathrm{cd}}(x)$ within both sets have measure zero, we show that the integral is well-defined. This is motivated by a classical result in real analysis: the integral of a bounded function is well-defined as a Riemann integral if and only if its set of discontinuities has Lebesgue measure zero.
For the set $\widetilde{\mathcal{X}}$, we use Assumption 2.1, which states that the bifurcation point set $\widetilde{\mathcal{X}}$ defined in (2.3) has measure zero. For the remaining set $\mathcal{X}\setminus\widetilde{\mathcal{X}}$, we make the following assumption:
@@
Under Assumption 2.1 and 3.4, $y^{\text{cd}}(x)$ is discontinuous only on a set of measure zero in $\mathcal{X}$. Since $\phi^{\text{cd}}(x)=f(x,y^{\text{cd}}(x))$ and the function $f$ is continuous, it follows that $\phi^{\text{cd}}(x)$ is also discontinuous only on a measure-zero subset of $\mathcal{X}$. Note that $\mathcal{X}$ is a closed convex set, and thus its boundary $\partial\mathcal{X}$ has Lebesgue measure zero (Lang, 1986, Theorem 1). The extension of $\phi^{\text{cd}}(x)$, i.e., setting $\phi^{\text{cd}}(x)=\overline{f}$ for $x\notin\mathcal{X}$, introduces new discontinuities only on a subset of $\partial\mathcal{X}$, since the function takes different values inside and outside $\mathcal{X}$. Therefore, the set of discontinuities of the extended $\phi^{\text{cd}}(x)$ on $\mathbb{R}^{n}$ remains a measure-zero set. By the classical criterion for Riemann integrability, it follows that $\phi^{\text{cd}}(x)$ is Riemann integrable on $\mathcal{X}$. Therefore, the second technical issue is resolved. In conclusion, the convolution $\phi^{\text{cd}}_{\xi}(x)$ is well-defined.
@@
Suppose Assumption 3.2 holds. Then $\phi^{\text{cd}}_{\xi}(x)$ is smooth for any $\xi$. Furthermore, if $\phi^{\text{cd}}(x)$ is continuous at point $\bar{x}$, then we have the convergence of the function value $\lim_{\xi\to 0}\phi^{\text{cd}}_{\xi}(\bar{x})=\phi^{\text{cd}}(\bar{x})$. If $\phi^{\text{cd}}(x)$ is differentiable at point $\bar{x}$, then we have the convergence of gradient $\lim_{\xi\to 0}\nabla_{x}\phi^{\text{cd}}_{\xi}(\bar{x})=\nabla_{x}\phi^{\text{cd}}(\bar{x})$, and thus $\lim_{\xi\to 0}P_{\mathcal{X}}(x,\nabla_{x}\phi^{\text{cd}}_{\xi},\beta)=P_{\mathcal{X}}(x,\nabla_{x}\phi^{\text{cd}},\beta)$ which are defined as follows:
@@
Proposition 3.2 shows that Gaussian smoothing gives $\phi^{\text{cd}}_{\xi}(x)$ good regularity: it is smooth for all $\xi>0$, and it recovers both the function value and (proximal) gradient of $\phi^{\text{cd}}$ in the limit as $\xi$ approaches 0, at any point where $\phi^{\text{cd}}$ is continuous or differentiable. This result ensures that the projected gradient-based methods applied to the smoothed problem can approximate the behavior of the original hyperfunction. The proof follows the classical theory of mollification in PDE (Evans, 2022), where a discontinuous or nonsmooth function is approximated by the convolution with a smooth kernel. In classical mollifier theory, a compactly supported smooth bump function (e.g., supported on the unit ball) is often used. Pointwise and gradient convergence are established using local continuity or differentiability inside the support, while the contribution from outside the support is exactly zero. In our case, the smoothing kernel is the Gaussian $h_{\xi}$, which is not compactly supported. Nevertheless, the key idea remains the same: most of the kernel mass is concentrated around the origin, and the contribution from the tail decays exponentially fast. This exponential decay plays the same role as compact support in standard mollifier arguments. Thus, the structure of the proof is essentially identical to the standard bump function case.
@@ 4.1 Sketch of the Proposed Algorithm
$=\frac{1}{\xi}\mathbb{E}_{u\sim\mathcal{N}(0,I_{n})}[uf(x+\xi u,y^{\text{cd}}(x+\xi u))].$
Note that, as established in (3.9), we have extended $\phi^{\text{cd}}$ outside $\mathcal{X}$, which is equivalently interpreted as redefining $f$ to take the constant value $\overline{f}$ outside $\mathcal{X}$, where $\overline{f}$ is from Proposition 3.1(2). Thus, in the following, we work with $f$ under this redefinition.
In practice, we estimate this gradient via Monte Carlo sampling, which leads to a biased stochastic gradient descent (SGD) algorithm. Specifically, we draw i.i.d. sample points $u^{(1)},...,u^{(N)}\sim\mathcal{N}(0,I_{n})$, and for each such sample $u^{(i)}$, if the perturbed point $x+\xi u^{(i)}\notin\mathcal{X}$, we directly assign the value $f(x+\xi u^{(i)},\hat{y}(x+\xi u^{(i)}))=\overline{f}$; otherwise, when $x+\xi u^{(i)}\in\mathcal{X}$, we run a fixed number $K$ of iterations of the lower-level algorithm $\mathcal{M}$ at $x+\xi u^{(i)}$ to obtain an approximate solution $\hat{y}(x+\xi u^{(i)})$, where $K$ is independent of $x$. We then form the following average
@@
         Run inner algorithm $\mathcal{M}$ on $g(\tilde{x}^{(i)},\cdot)$ for $K_{t}$ steps, initialized at $y_{0}$, with step size $\bm{\eta}$ to obtain $\hat{y}^{(i)}$;
@@
  Compute the gradient estimator $\hat{\nabla}_{x}^{K_{t}}\phi^{\text{cd}}_{\xi}(x_{t})$ given by (4.3) and update $x_{t+1}$: $\hat{\nabla}_{x}^{K_{t}}\phi^{\text{cd}}_{\xi}(x_{t})\leftarrow\frac{1}{N_{t}\,\xi}\sum_{i=1}^{N_{t}}u^{(i)}f^{(i)}$;
@@
$\nabla_{x}\phi^{\text{cd}}_{\xi}(x)-\nabla^{K}_{x}\phi^{\text{cd}}_{\xi}(x)=\int_{\mathbb{R}^{n}}\nabla_{x}h_{\xi}(x-z)[f(z,y^{\text{cd}}(z))-f(z,\hat{y}(z))]dz.$ (4.5)
@@
$\int_{\widetilde{\mathcal{X}}_{\delta}}\nabla_{x}h_{\xi}(x-z)[f(z,y^{\text{cd}}(z))-f(z,\hat{y}(z))]\,dz+\int_{\mathbb{R}^{n}\setminus\widetilde{\mathcal{X}}_{\delta}}\nabla_{x}h_{\xi}(x-z)[f(z,y^{\text{cd}}(z))-f(z,\hat{y}(z))]\,dz.$ (4.7)
@@
Finally, to obtain the lower-level problem oracle complexity of SCiNBiO, we introduce the notion of fold bifurcation Section 4.4 to endow the bifurcation points with additional structure. This also leads to a clearer geometric picture of the bifurcation set.